# Evidence of de Almeida–Thouless line below six dimensions

### v1

@@ 1 Introduction
Some of the most basic notions in the theory of critical phenomena and the renormalization group (RG) [1, 2, 3] become suspect when the magnetically disordered alloys known as spin glasses [4, 5, 6] are subjected to an externally applied magnetic field. We are referring, in particular, to the lower $d_{\text{l}}$ and the upper critical dimension $d_{\text{u}}$ —no phase transition appears if the space dimension is $d<d_{\text{l}}$, while Mean Field approximation provides correct predictions for the critical exponents if $d>d_{\text{u}}$. Indeed, the value of $d_{\text{l}}$ and $d_{\text{u}}$ still remains controversial in spite of the significant uninterrupted efforts of the community [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. This confusing scenario is somewhat surprising given the extraordinary success of the RG in the study of disordered systems, including classical [32, 33] and quantum [34, 35] spin glasses in zero magnetic field.
@@
Non perturbative methods, such as Monte Carlo simulations, could be very useful to shed light although, as it has been recently emphasized [25, 26], they suffer from other limitations such as the limited range of system sizes that can be equilibrated. The quasi-first order phase transition scenario has been considered recently [37, 31], without finding any evidence supporting it. The $d_{\text{l}}=\infty$ seems to have been ruled out as well. Indeed, numerical evidence for a dAT line in $d=6$ has been found [31]. Also, a high-temperature series finds a dAT line for $d\geq 6$ [16]. The finding of a dAT line exactly at $d=6$ (but not for $d>6$) has been disputed by Kac-model based studies [23, 24, 25, 26] —the Kac-model is a one-dimensional model with long range interactions that mimics the behavior of systems with short-range interactions in $d$ spatial dimensions— but other studies claim in favor of a dAT line for much lower $d$ [27, 28]. According to Refs. [25, 26], the finding of a dAT line at $d=6$ would have been an artifact caused by the small size of the systems that could be brought to thermal equilibrium, but the existence of a dAT line for $d>6$ is undisputed.
@@ 5.1 Critical temperature and critical exponents
Finally, using the values obtained for $\omega=3.7(4)$ and $\nu=0.729(11)[2]$ we can extrapolate to an infinite system size of the critical temperature
@@ 5.2 Lower critical dimension
With those values of the exponents, we try several fits for $1/\nu(d)$ and $\eta(d)-2+d$. We find that a linear function does not describe the behavior of the critical exponents. For $1/\nu(d)$, we choose to fit to a quadratic function $f^{\nu}_{\mathrm{fit}}(d)=2+b(6-d)+c(6-d)^{2}$ that gives $\chi^{2}/\mathrm{d.o.f}=0.0028$ and a $p$-value of $95\%$. The function $f^{\nu}_{\mathrm{fit}}$ is chosen to impose that at the upper critical dimension $d_{U}=6$ we recover the Gaussian value of $\nu=1/2$. For $\eta$, we prefer a third degree fit of $\eta(d)-2+d$ to $f^{\eta}_{\mathrm{fit}}(d)=4+b_{1}(6-d)+b_{2}(6-d)^{2}+b_{3}(6-d)^{3}$. Again, we choose the function so that for $d=6$ we recover $\eta(d)-2+d=4$, i.e., the MF value $\eta=0$.
@@
Figure 3: Critical exponents $\eta$ and $\nu$ as a function of the dimension for $h=0$. The points correspond to the value of the critical exponents that appear in Table 5. The solid lines correspond to the fits to the functions $f^{\nu}_{\mathrm{fit}}$ and $f^{\eta}_{\mathrm{fit}}$ defined in the text. The horizontal dashed lines at zero are used for reference to the eye.
@@ 7 Conclusions
In spite of the above caveats, we do have found clear numerical evidence for a de Almeida-Thouless line in $d=5$. We have also been able to compute the critical exponents, including the leading corrections to scaling exponent. Of course, as in any other numerical investigation, one can never discard that qualitatively new behavior will eventually appear when still larger systems will be equilibrated (for more on this viewpoint, see Refs. [25, 26]). Yet, out findings in $d=5$ are consistent with previous results obtained in four dimensions [20], as well as in the Kac model [27].
@@ 8 Acknowledgments
We thank M. C. Angelini, S. Palazzi, G. Parisi, and T. Rizzo for interesting discussion and sharing the results of their analytical computation. We acknowledge partial financial support from Grant Nos. PID2024-156352NB-I00, PID2022-136374NB-C21, PID2020-112936GB-I00 funded by MCIN/AEI/10.13039/501100011033, and from Grant No. GR24022 funded by Junta de Extremadura (Spain) and by European Regional Development Fund (ERDF) “A way of making Europe. We have performed the numerical simulations in the computing facilities of the Instituto de Computación Científica Avanzada (ICCAEx).

### v3

@@ 1 Introduction
Some of the most basic notions in the theory of critical phenomena and the renormalization group (RG) [1, 2, 3] become suspect when the magnetically disordered alloys known as spin glasses [4, 5, 6] are subjected to an externally applied magnetic field. We are referring, in particular, to the lower $d_{\text{l}}$ and the upper critical dimension $d_{\text{u}}$ —no phase transition appears if the space dimension is $d<d_{\text{l}}$, while Mean Field approximation provides correct predictions for the critical exponents if $d>d_{\text{u}}$. Indeed, the value of $d_{\text{l}}$ and $d_{\text{u}}$ still remains controversial in spite of the significant uninterrupted efforts of the community [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. This confusing scenario is somewhat surprising given the extraordinary success of the RG in the study of disordered systems, including classical [33, 34] and quantum [35, 36] spin glasses in zero magnetic field.
@@
Non perturbative methods, such as Monte Carlo simulations, could be very useful to shed light although, as it has been recently emphasized [25, 26, 27], they suffer from other limitations such as the limited range of system sizes that can be equilibrated. The quasi-first order phase transition scenario has been considered recently [38, 32], without finding any evidence supporting it. The $d_{\text{l}}=\infty$ seems to have been ruled out as well. Indeed, numerical evidence for a dAT line in $d=6$ has been found [32]. Also, a high-temperature series finds a dAT line for $d\geq 6$ [16]. The finding of a dAT line exactly at $d=6$ (but not for $d>6$) has been disputed by Kac-model based studies [23, 24, 25, 26, 27] —the Kac-model is a one-dimensional model with long range interactions that mimics the behavior of systems with short-range interactions in $d$ spatial dimensions— but other studies claim in favor of a dAT line for much lower $d$ [28, 29]. According to Refs. [25, 26], the finding of a dAT line at $d=6$ would have been an artifact caused by the small size of the systems that could be brought to thermal equilibrium, but the existence of a dAT line for $d>6$ is undisputed.
@@ 5.1 Critical temperature and critical exponents
Finally, using the values obtained for $\omega=3.7(4)$ and $\nu=0.729(11)[2]$ we can extrapolate to an infinite system size of the critical temperature  [^3] We have verified that our estimate of $\beta_{c}^{\infty}$ is stable under the inclusion of a second correction term in the fitting form, $\beta_{c}(L_{1})=\beta_{c}^{\infty}+A\,L_{1}^{-\omega-1/\nu}+B\,L_{1}^{-2\omega-1/\nu}$. The resulting estimates are fully consistent with those obtained from the one–correction fit.
@@ 5.2 Lower critical dimension
With those values of the exponents, we try several fits for $1/\nu(d)$ and $\eta(d)-2+d$. We find that a linear function does not describe the behavior of the critical exponents. For $1/\nu(d)$, we choose to fit to a quadratic function $f^{\nu}_{\mathrm{fit}}(d)=2+b(6-d)+c(6-d)^{2}$ that gives $\chi^{2}/\mathrm{d.o.f}=0.0028$ and a $p$-value of $95\%$. The function $f^{\nu}_{\mathrm{fit}}$ is chosen to impose that at the upper critical dimension $d_{U}=6$ we recover the Gaussian value of $\nu=1/2$. For $\eta$, we prefer a third degree fit of $\eta(d)-2+d$ to $f^{\eta}_{\mathrm{fit}}(d)=4+b_{1}(6-d)+b_{2}(6-d)^{2}+b_{3}(6-d)^{3}$. Again, we choose the function so that for $d=6$ we recover $\eta(d)-2+d=4$, i.e., the MF value $\eta=0$  [^4] Because we have numerical data in $d=3,4,5$ and impose the analytic constraint $\eta(6)=0$ at the upper critical dimension, a cubic fit contains three free parameters for three numerical points, leaving zero residual degrees of freedom. Consequently, no standard estimate of the fit uncertainty can be provided. The uncertainty in the crossing point $d_{L}$ is instead obtained through a parametric bootstrap (Gaussian resampling): each data point is perturbed according to its numerical error, $10^{3}$ synthetic datasets are generated, and the standard deviation of the resulting $d_{L}$ values is taken as the error bar..
@@
Figure 3: Critical exponents $\eta$ and $\nu$ as a function of the dimension for $h=0$. The points correspond to the value of the critical exponents that appear in Table 5. The solid lines correspond to the fits to the functions $f^{\nu}_{\mathrm{fit}}$ and $f^{\eta}_{\mathrm{fit}}$ defined in the text. The horizontal dashed lines at zero are used for reference to the eye. Statistical errors are present but, in most cases, smaller than the symbol size.
@@ 7 Conclusions
In spite of the above caveats, we do have found clear numerical evidence for a de Almeida-Thouless line in $d=5$. We have also been able to compute the critical exponents, including the leading corrections to scaling exponent. Of course, as in any other numerical investigation, one can never discard that qualitatively new behavior will eventually appear when still larger systems will be equilibrated (for more on this viewpoint, see Refs. [25, 26, 27]). Yet, out findings in $d=5$ are consistent with previous results obtained in four dimensions [20], as well as in the Kac model [28].
@@ 8 Acknowledgments
We thank M. C. Angelini, S. Palazzi, G. Parisi, and T. Rizzo for interesting discussion and sharing the results of their analytical computation. We also thank Giorgio Parisi for discussions about our numerical results. We acknowledge partial financial support from Grant Nos. PID2024-156352NB-I00, PID2022-136374NB-C21, PID2020-112936GB-I00 funded by MCIN/AEI/10.13039/501100011033/FEDER, UE. and from Grant No. GR24022 funded by Junta de Extremadura (Spain) and by European Regional Development Fund (ERDF) “A way of making Europe. We have performed the numerical simulations in the computing facilities of the Instituto de Computación Científica Avanzada (ICCAEx).